Dr Paul Coddington NPAC, Rm 1-019 CST 111 College Place Syracuse University Syracuse, NY 13244

Tel: (315) 443-4883 FAX: (315) 443-1937

Email: paulc@npac.syr.edu

12/3/92

Dear Ms Gagnon,

I would like to request a copy of the EPRI EMTP software. I am leading a research

project at the Northeast Parallel Architectures Center at Syracuse University investigating the use of parallel processing for power systems applications. In particular we are studying parallel algorithms for sparse matrix and differential-algebraic equation solvers. We would like to test our algorithms on real power utility software, and would appreciate the use of your software for this project.

Yours sincerely,